OpenStudy (anonymous):

The ratio of two numbers is 7 to 5. The sum of the numbers is 36. What are the numbers?

OpenStudy (anonymous):

let the two numbers be 7x and 5x 7x + 5x = 36 12x = 36 dividing both sides by 12 x = 3 two numbers are 7*3 = 21 5*3 = 15

OpenStudy (turingtest):

Sheg's answer is more efficient than mine, but here's another way to see it: "The ratio of two numbers is 7 to 5" translated into math language is\[{x \over y}={7\over5}\]"The sum of the numbers is 36" translates as\[x+y=36\]2 equations+2 unknowns=A solvable problem! Sheg has provided the answer :)